4. Find local extremizers for the following optimization problem Minimize f(x1,x2)=xi + 22, subject to 3r+4z12 +62 140.

4. Find local extremizers for the following optimization problem:
Answer #1


To find the minimum values of/倆両)-而2 +ろ2 subject to 3x12 +4x1ろ+ 6x22 = 140 Thelagrangean function is 2x +2(6 +4%)=0 6x1 +4x2

Case-2 %,-14 ence x1 12 Hence the points of extrema are (2,4 = 4(14) +14 = 70 (12,14) = (-2)2 +(H)2 = 4+16 = 20 (±2,14) = 20